Write down your assigned network address: 224.151.41.0/24

1. What is your assigned network address?

2. What is the subnet mask (dotted decimal form) for this network?

3. What is the broadcast address for this network?

4. How many addressable hosts are available for this network?

Following good network design practices, you decide to segment this network into 4 equal subnets, to be used or shared among the different Rowdy Gamerz departments: Software Development, Technical Support, Marketing/Sales, and Management & Human Resources.

5. To create 4 subnets, you must borrow how many bits from the host portion of the network? (Hint: solve 4 = 2n)

6. What is the subnet mask (dotted decimal form) for these new subnets?
Answer #1

1. Assigned network address: 224.151.41.0 / 24

2. Subnet mask: 255.255.255.0 ( /24 )

3. Broadcast address: Wild Card Mask + Network Address

Subnet mask => 11111111.11111111.11111111.00000000

Wild Card Mask => 00000000.00000000.00000000.11111111

Network Address => 11100000.10010111.00101001.00000000

So, result of OR operation between wild card mask and network address is:

11100000.10010111.00101001.11111111

i.e., Broadcast address is: 224.151.41.255

4.

Host bits = 32 - 24 = 8

Number of hosts = = 256

Number of addressable hosts = 256 - 2 = 254

5.

To create 4 subnets, we need

4 =

i.e., we need to borrow 2 bits from host bits portion.

6.

We have borrowed 2 bits from host into network portion: 24 + 2 = 26

i.e., 26 bits for Network and 6 bits for Host

So, New subnet mask is: 255.255.255.192